OpenStudy (experimentx):

For \( a, b, c > 0, abc =1\) show that \[ \frac{a}{b}+\frac{b}{c}+\frac{c}{a}\ge a+b+c \] No Lagrange multiplier allowed.

OpenStudy (anonymous):

*

OpenStudy (anonymous):

\[a^2c+ab^2+bc^2\ge a+b+c\]

OpenStudy (experimentx):

another form would be it. or \[ \frac{ab^2+bc^2+ca^2}{a+b+c} \ge 1\]

OpenStudy (experimentx):

most likely this problem won't take more than AM-GM

OpenStudy (anonymous):

i want to say the next step is reducing it to 2 variables but i dont want to think, good luck to the next person

OpenStudy (experimentx):

it took nearly week for me to figure out it's solution. try using this technique http://en.wikipedia.org/wiki/Inequality_of_arithmetic_and_geometric_means#Weighted_AM.E2.80.93GM_inequality it will be lot shorter.